# Exploring the AGN Fraction of a Sample of JWST’s Little Red Dots at : Overmassive Black Holes Are Strongly Favored

### v1

@@ 5 Discussion and Conclusions
However, despite these intrinsic limitations, our work has shown that rigorously probing the parameter space in the stellar and black hole mass can constrain our sample of LRDs as high- AGN hosts, as also suggested by previous studies (Greene et al., 2024) of different samples. In our sample of eight LRDs, we found that the best fit AGN contribution is between 40% to 85%. The inferred AGN bolometric luminosities show black holes that are significantly overmassive compared to the stellar mass of the host, based on expectations from the local Reines & Volonteri (2015) . These sources are in better agreement with the high-z relation inferred by Pacucci et al. (2023) from JWST galaxies in the redshift range .
@@
The fact that most of the black holes observed thus far by JWST at are overmassive with respect to the host’s stellar content is not necessarily surprising. Earlier works (e.g., Volonteri et al. 2023) predicted that if JWST were to observe central black holes in high-, smaller galaxies, they had to be overmassive. However, Pacucci et al. (2023) recently found that the entire dataset of galactic systems found thus far by JWST at is so overmassive, and JWST’s sensitivity in H so deep, that their ratios cannot be explained by being extreme outliers from the local relation. They seem to be, in fact, a different statistical population of objects, with black hole masses significantly overmassive. Intriguingly, these galactic systems seem to be located close to the local and relations, suggesting that these two relationships are more fundamental than the (Maiolino et al., 2023; Juodžbalis et al., 2024). This population of sources is also X-ray weak, with several studies cross-matching samples of LRDs with X-ray catalogs and finding no or very few detections (see, e.g., Ananna et al. 2024; Yue et al. 2024a; Maiolino et al. 2024b). One possibility to solve this tension would be to assume that the IR emission is overestimating the AGN luminosity, leading to lower black hole masses. Unusually large levels of obscuration could also play a part in explaining the deficiency of X-ray emissions.
@@
Acknowledgments: We thank Guillermo Barro for providing the photometry of the eight sources investigated here. E.D. acknowledges support from a Dartmouth Fellowship. F.P. thanks Dale Kocevski for clarifying the differences in the selection of EROs and LRDs; our “Note on Terminology” (Sec. 2.1) is based on his insights. F.P. acknowledges support from a Clay Fellowship administered by the Smithsonian Astrophysical Observatory. R.C.H. acknowledges support from NASA through Astrophysics Data Analysis Program grant 80NSSC23K0485. This work was also supported by the Black Hole Initiative at Harvard University, which is funded by grants from the John Templeton Foundation and the Gordon and Betty Moore Foundation.

### v2

@@ 5 Discussion and Conclusions
However, despite these intrinsic limitations, our work has shown that rigorously probing the AGN vs stellar contribution to the SED can constrain our sample of LRDs as high- AGN hosts, as also suggested by previous studies (Greene et al., 2024) of different samples. In our sample of 19 LRDs, we generally found that the best fit AGN contribution is between 40% to 85%. The inferred AGN bolometric luminosities show black holes that are significantly overmassive compared to the stellar mass of the host, based on expectations from the local Reines & Volonteri (2015) . These sources better agree with the high-z relation inferred by Pacucci et al. (2023) from JWST galaxies in the redshift range .
@@
The fact that most of the black holes observed thus far by JWST at are overmassive with respect to the host’s stellar content is not necessarily surprising. Earlier works (e.g., Volonteri et al. 2023) predicted that if JWST were to observe central black holes in high-, smaller galaxies, they had to be overmassive. However, Pacucci et al. (2023) recently found that the entire dataset of galactic systems found thus far by JWST at is so overmassive, and JWST’s sensitivity in H so deep, that their ratios cannot be explained by being extreme outliers from the local relation. They seem to be, in fact, a different statistical population of objects, with black hole masses significantly overmassive. Intriguingly, these galactic systems seem to be located close to the local and relations, suggesting that these two relationships are more fundamental than the (Maiolino et al., 2023; Juodžbalis et al., 2024). This population of sources is also X-ray weak, with several studies cross-matching samples of LRDs with X-ray catalogs and finding no or very few detections (see, e.g., Ananna et al. 2024; Yue et al. 2024a; Maiolino et al. 2024b). One possibility to solve this tension would be to assume that the IR emission is overestimating the AGN luminosity, leading to lower black hole masses. Unusually large levels of obscuration (Maiolino et al., 2024b) or mildly super-Eddington accretion levels, leading to intrinsically X-ray weak SEDs (Pacucci & Narayan, 2024), could also explain the X-ray deficiency; the X-ray faintness of the LRDs would then require next-generation, high-resolution X-ray observatories, such as AXIS (Reynolds et al., 2023).
@@
Acknowledgments: We thank Pablo Perez-Gonzalez for providing the photometry of the 19 sources investigated here. E.D. acknowledges support from a Dartmouth Fellowship. F.P. acknowledges support from a Clay Fellowship administered by the Smithsonian Astrophysical Observatory. R.C.H. acknowledges support from NASA through Astrophysics Data Analysis Program grant 80NSSC23K0485. This work was also supported by the Black Hole Initiative at Harvard University, which is funded by grants from the John Templeton Foundation and the Gordon and Betty Moore Foundation.